If we write the following complex number in standard form:
(sqrt7 + sqrt11 i)(sqrt7 - sqrt11 i) = a + bi
Then a = ? and b = ?

Answer by stanbon(48545)

If we write the following complex number in standard form:
(sqrt7 + sqrt11 i)(sqrt7 - sqrt11 i) = a + bi
= 7 - 11i^2
= 7 + 11
= 18 + 0i
Then a = 18 and b = 0
---------------------------
Cheers,
Stan H.